# ReVision: High-Quality, Low-Cost Video Generation with Explicit 3D Physics Modeling for Complex Motion and Interaction

### v1

@@ 6.2 Complex Motion Generation
Experimental Setup. To demonstrate our model’s ability to generate videos with complex motion, we compare our approach with state-of-the-art human image animation models on the TikTok Dancing dataset [19], using the Disco [61] split. For compatibility with the SVD model architecture, all videos are cropped to $576\times 1024$. We fine-tune the original SVD only on the training split for 30K iterations, with a batch size of 8 and a learning rate of $1\times 10^{-5}$.
Evaluation Metrics. We follow baselines and report Peak Signal-to-Noise Ratio (PSNR) [15], Structural Similarity Index (SSIM) [63], and Learned Perceptual Image Patch Similarity (LPIPS) [76] to measure the visual quality of the generated results. We also report and Fréchet Video Distance (FVD) [56] for video fidelity comparision.
Experimental Results. We compare ReVision with human image animation methods in Tab. 2, where our approach achieves state-of-the-art performance across all metrics. Notably, we observe a significant improvement in FVD, highlighting substantial gains in video generation quality. It is important to note that all baseline methods in this task rely on ground truth motion sequences, which are challenging to obtain in practical scenarios, thus limiting their applicability. In contrast, our method can generate realistic and high-quality videos using only the inference image and the target pose as input.
@@
Disco [61] 0.668 29.03 0.292 292.80
MagicAnimate [67] 0.714 29.16 0.239 179.07
Animate Anyone [16] 0.718 29.56 0.285 171.90
Champ [78] 0.802 29.91 0.234 160.82
VividPose [60] 0.758 29.83 0.261 152.97
ReVision (w/ full) 0.842 30.05 0.227 137.53
ReVision - - - 130.14
Table 2: Quantitative comparison for human dance generation. ‘ReVision (w/ full)’ follows all baselines and takes the full motion sequences as condition, while ‘ReVision’ generates videos with only the target pose from the final frame.
6.3 Ablation Study
Herein, we conduct ablation studies to verify the effectiveness of the proposed designs.
Figure 6: The parametric 3D mesh serves as an effective object-level prior, ensuring complete human body structures in the coarse video generated during the first stage. In the left two images, the human keypoint model fails to detect the missing right hand, which is accurately “recovered” by the parametric human mesh model. In the right two images, the human mesh model provides a more accurate prior for both blurred hands.
Parametric 3D Mesh. Previous human image animation models mainly rely on 2D pose sequences for each frame to provide motion information. However, this approach is not optimal for general video generation. As shown in Fig. 6, we compare the results of using a parametric human mesh model [32] versus a human keypoint model [54]. Our findings indicate that the human mesh model provides a robust object-level prior, which significantly benefits general video generation. Specifically, current video generation models often misinterpret the structure of humans and animals, occasionally producing unrealistic results, such as a man with three arms, an example of morphological failure. This problem becomes more serious in complex motion generation. However, incorporating human and animal priors from 3D mesh models substantially mitigates these structural inaccuracies, enabling more accurate representations of targets.
Object Consistency $\uparrow$ Motion Consistency $\uparrow$ Morphological Failure Rate $\downarrow$
Stage 1 12.4 4.0 83.5
Stage 3 87.6 96.0 14.3
Table 3: User preference studies for multi-stage generation. By explicitly using the 3D knowledge during generation, videos generated in stage 3 show significant improvements in object and motion consistency, while reducing the morphological failure rate.
@@ 6.3 Ablation Study
Multi-Stage Generation. To demonstrate the effectiveness of multi-stage generation, we select a complex dance scenario and visualize outputs at different stages in Fig. 7: the coarse video from stage one, the optimized 3D mesh from stage two, and the re-generated video from stage three. We also present quantitative improvements of our multi-stage generation in Tab. 3, where 500 video pairs were evaluated by users on Amazon MTurk. Each pair was rated by three users, resulting in a total of 1,500 evaluations. The results show that the video generation model alone still struggles to produce realistic videos with accurate motion. However, leveraging object-level priors from the parametric 3D mesh model and motion priors from our PPPM enables the generation of a realistic 3D motion sequence aligned with the reference image, target pose, text prompt, and motion strength. Guided by this 3D motion sequence, Stage 3 generates a realistic video with enhanced motion and visual quality.
@@
In addition, Fig. 7 shows that even when the first stage generates videos with significantly broken motions, the second stage can recover a relatively smooth and reasonable motion sequence using the accurate first frame and target pose, ultimately ensuring successful final video generation. This demonstrates the robustness of our pipeline.
Figure 7: Video generated in stage 1 (S1) and stage 3 (S3). Although fine-tuned on a dataset containing large motions, SVD still struggles to generate complex motion sequences. However, by incorporating the optimized 3D dense motion sequence from Stage 2, we significantly enhance both motion and visual quality and generate realistic videos with large motions in stage 3.
7 Limitations
The proposed method has several remaining limitations. First, it relies on parametric 3D mesh models, requiring multiple off-the-shelf models for different object categories, though it adds only 5 seconds to the total inference time. Recent advances in 3D modeling, such as encoding 3D priors of general objects within a single diffusion model [30], are paving the way for more general, efficient models that can be seamlessly integrated into our pipeline for high-quality video generation. Second, the model still struggles to generate high-quality details such as fingers and hands. Finally, while PPPM can generate realistic motion sequences beyond 32 frames, our current implementation, based on vanilla SVD, is limited by memory constraints (80GB RAM). However, recent methods have demonstrated longer video generation ability using pretrained diffusion models [9]. Exploring long-video generation with 3D knowledge remains future work.
@@ 8 Conclusion
In this work, we introduced ReVision, a three-stage framework for video generation that improves motion consistency by integrating 3D motion cues. Our approach leverages a pretrained video diffusion model to generate coarse videos, refines 3D motion sequences through PPPM, and reconditions the generation process with these enhanced motions to improve fine-grained and complex motion generations. Evaluations show that ReVision significantly outperforms existing methods in motion fidelity and coherence.
Supplementary Material
@@
  • Architectural details for PPPM. (Sec. 9)
  • Quantitative ablations of multi-stage generation (Sec. 10)
@@
9 Architectural details for PPPM
@@
To optimize the 3D motion sequence extracted from the coarse generated video, we propose the Parameterized Physical Prior Model (PPPM). As shown in Fig. 8, PPPM utilizes a transformer architecture with self-attention, cross-attention, and feedforward layers as its backbone. It takes the parameterized motion sequence of the coarse video as input and optimizes it based on the input text prompt and motion strength.
10 Improvement of Multi-stage Generation
@@
We briefly discuss the quantitative ablations of multi-stage generation in the main paper and provide additional experimental details and results here. To further demonstrate the improvements of our multi-stage approach, we conducted an additional user study on Amazon MTurk comparing videos from Stages 1 and 3. Unlike our previous study, which compared our method with SVD, this evaluation focuses on object consistency and motion consistency. We also report the percentage of videos containing incorrect human or animal structures (i.e., the morphological failure rate). We evaluate 500 video pairs, each rated by three different users, resulting in 1,500 total evaluations. The results are presented in Tab. 3 of the main paper. By optimizing with a parametric 3D mesh, our approach significantly reduces incorrect human and animal structures, leading to substantial improvements in object and motion consistency.
Figure 8: Architecture of the Parameterized Physical Prior Model

### v2

@@ 5.2 Complex Motion Generation
Experimental Setup. To demonstrate our model’s ability to generate videos with complex motion, we compare our approach with state-of-the-art human image animation models on the TikTok Dancing dataset (Jafarian & Park, 2021), using the Disco (Wang et al., 2024c) split. For compatibility with the SVD model architecture, all videos are cropped to $576\times 1024$. We fine-tune the original SVD only on the training split for 30K iterations, with a batch size of 8 and a learning rate of $1\times 10^{-5}$.
Figure 8: PMP improves motion and visual quality, generating realistic videos with large motions.
Evaluation Metrics. We follow baselines and report Peak Signal-to-Noise Ratio (PSNR) (Hore & Ziou, 2010), Structural Similarity Index (SSIM) (Wang et al., 2004), and Learned Perceptual Image Patch Similarity (LPIPS) (Zhang et al., 2018) to measure the visual quality of the generated results. We also report and Fréchet Video Distance (FVD) (Unterthiner et al., 2018) for video fidelity comparision.
Experimental Results. We compare ReVision with human image animation methods in Tab. 6, where we achieve state-of-the-art performance across all metrics. Notably, we observe a significant improvement in FVD, highlighting substantial gains in video generation quality. It is important to note that all baselines in this task rely on ground-truth motion sequences, which are challenging to obtain in practical scenarios, limiting their applicability. In contrast, our method can generate realistic and high-quality videos using only the input inference image or inference image with a target pose.
Table 5: Quantitative comparisons for dance generation. ‘ReVision (w/ full motion)’ follows baselines and takes full motion sequences as condition, while ‘ReVision (w/ target pose)’ uses the target pose from the final frame.
@@
MagicAnimate (Xu et al., 2024) 0.714 29.16 0.239 179.07
Animate Anyone (Hu, 2024) 0.718 29.56 0.285 171.90
Champ (Zhu et al., 2024) 0.802 29.91 0.234 160.82
VividPose (Wang et al., 2024b) 0.758 29.83 0.261 152.97
ReVision (image only) - - - 136.43
ReVision (w/ target pose) - - - 130.14
ReVision (w/ full motion) 0.864 30.08 0.210 121.26
Table 6: PMP acts as a general motion denoiser, improving performance on human motion generation. Following MoMask, we report R-Precision at Top-1, Top-2, and Top-3. Our PMP achieves state-of-the-art performance on two widely used benchmarks.
HumanML3D KIT-ML
R-P@1 R-P@2 R-P@3 R-P@1 R-P@2 R-P@3
MotionDiffuse (Zhang et al., 2024b) 0.491 0.681 0.782 0.417 0.621 0.739
ReMoDiffuse (Zhang et al., 2023b) 0.510 0.698 0.795 0.427 0.641 0.765
MoMask (Guo et al., 2024) 0.521 0.713 0.807 0.433 0.656 0.781
MoMask + PMP 0.544 0.735 0.810 0.471 0.673 0.785
5.3 Parameterized Motion Prior Model (PMP)
We demonstrate the effectiveness of the proposed Parameterized Motion Prior model in this section.
Table 7: User studies for PMP. PMP improves object and motion consistency, while reducing morphological failure rates.
Object Motion Morphological
Consistency $\uparrow$ Consistency $\uparrow$ Failure Rate $\downarrow$
w/o PMP 12.4 4.0 83.5
w/ PMP 87.6 96.0 14.3
@@ 5.3 Parameterized Motion Prior Model (PMP)
PMP Enables High-Quality Video Generation with Complex Motions and Interactions. To demonstrate the effectiveness of PMP, we select a complex dance scenario and visualize outputs with and without the proposed PMP in Fig. 8. We also show quantitative improvements of the generated videos with PMP in Tab. 7, where 500 video pairs were evaluated by random users on Amazon MTurk. Each pair was rated by three different users, resulting in a total of 1,500 evaluations. The results show that the video generation model alone still struggles to produce high-quality videos with accurate motion. However, leveraging the object-level priors from our Parameterized Motion Prior enables the generation of realistic videos with enhanced motion and visual quality.
@@
PMP Prevents Error Accumulation in Multi-stage Video Generation. In addition, Fig. 8 shows that even when the generated videos exhibit severely broken motion, our PMP can still recover (predict) a smooth and coherent motion sequence using the ground-truth first frame and target pose, enabling successful final video generation. This correction mitigates motion errors and prevents error accumulation, highlighting the robustness of our pipeline.
PMP as a General Motion Denoiser. We focus on a more specific human motion generation task and show that our model improves the performance of the state-of-the-art method, MoMask (Guo et al., 2024), on standard benchmarks (see Tab. 6). Specifically, we use PMP to refine the motion sequences generated by MoMask and compare the results with MoMask and other methods on HumanML3D (Guo et al., 2022) and KIT-ML (Plappert et al., 2016) benchmarks. We adopt the same training dataset as MoMask and apply the perturbations described in Sec. 4.2 to train PMP. Serving as a general motion denoiser, our model consistently enhances motion generation quality of the current best models across multiple benchmarks.
@@ 6 Conclusion
We introduced ReVision, a three-stage framework for video generation that improves motion consistency by integrating 3D motion cues. ReVision leverages a pretrained video diffusion model to generate coarse videos, refines 3D motion sequences with PMP, and reconditions the generation process with enhanced motions to improve fine-grained and complex motion generations. Evaluations show that ReVision significantly outperforms existing methods in motion fidelity and coherence.
Acknowledgements. Thanks to all who supported this project and to the anonymous reviewers for their constructive comments. AY acknowledges support from the ONR N000142412696.
Appendix
@@
  • Sec. A provides the architectural details for PMP.
  • Sec. B provides additional details for our annotated dataset.
  • Sec. C provides additional ablation studies omitted from the main paper.
  • Sec. D discusses the limitations of our method.
  • Sec. E discusses the societal impacts of our method.
@@
Appendix A Architectural details for PMP
Figure 9: Architecture of the Parameterized Motion Prior model
@@ Appendix A Architectural details for PMP
To optimize the 3D motion sequence extracted from the coarse generated video, we propose the Parameterized Motion Prior model (PMP). As shown in Fig. 9, PMP utilizes a transformer architecture with self-attention, cross-attention, and feedforward layers as its backbone. It takes the parameterized motion sequence of the coarse video as input and optimizes it based on the input text prompt and motion strength. More specifically, PMP is trained as a single shared model across all categories. It directly predicts the corrected pose from the perturbed input, and we supervise it with an MSE loss between the predicted pose and the ground-truth pose.
Appendix B Dataset
Both the motion-conditioned video generation model and the Parameterized Motion Prior model (PMP) need to be fine-tuned (trained) on a small yet high-quality video dataset with object-centric annotations. Existing large-scale video datasets (Bain et al., 2021; Chen et al., 2024) mainly provide text-image pairs without detailed object-centric annotations. To address this limitation, we use a suite of off-the-shelf models across various tasks to generate detailed 2D and 3D object-centric annotations. We annotate a total of 20K videos from the Panda-70M (Chen et al., 2024) dataset, with approximately 55% human videos, 15% animal videos, and 30% general-object videos. For each video, we provide frame-wise 2D bounding boxes, semantic masks, depth estimation maps, and 3D parametric mesh reconstructions for detected humans and animals. The details are outlined below.
High-Quality Motion Videos Filtering. To start with, we use LLMs (Yang et al., 2024a) and an open-vocabulary segmentation model (Yu et al., 2023) to curate high-quality motion videos. Specifically, LLM filters videos with evident motion based on their captions. Then, for each selected video, we equally sample 10 frames and apply the segmentation model to identify humans and animals. We evaluate each frame based on the predicted mask size and mask count. Then we retain videos where humans or animals occupy a significant portion of the frame and where the count of humans does not exceed five in each frame.
Object Detection and Depth Estimation. Based on the captions of videos, we identify the objects mentioned and detect their bounding boxes (Varghese & Sambath, 2024) and instance masks (Yu et al., 2023). We also apply Depth Anything V2 (Yang et al., 2024b) to generate the depth maps of each frame.
Human Videos Annotation. For videos containing humans, we focus on extracting 2D instance segmentation masks, 2D part masks, 2D face keypoints, 3D body pose and shape, and 3D hand pose. We begin by using YOLO-V8 (Varghese & Sambath, 2024) to segment all humans in each frame, providing accurate human masks. Next, we apply a state-of-the-art face keypoint detector, RTMPose (Jiang et al., 2023), to predict facial keypoints for each detected human. Simultaneously, we use 4D-Human (Goel et al., 2023) and HaMeR (Pavlakos et al., 2024) to estimate the 3D body and hand meshes. The resulting SMPL (body mesh) (Loper et al., 2015) and MANO (hand mesh) (Romero et al., 2022) parameters are then fit into a unified SMPL-X (Pavlakos et al., 2019) representation, which contains both human body and hand meshes. We then project the 3D SMPL-X human mesh onto 2D to obtain part masks, as each vertex in the SMPL-X mesh is labeled by body part. Finally, we project the face keypoints and 3D human mesh onto the instance mask, allowing us to compute the overlap between the projected keypoints, projected human mask, and detected 2D human mask. This overlap is quantified using an IoU score, which is used to filter out annotations with high errors. As a result, for each video, we obtain annotations including human instance masks, 2D facial keypoints, 3D SMPL-X meshes for the body and hands, and 2D part-level segmentation masks.
Animal Videos Annotation. We start by using Grounded SAM 2 (Kirillov et al., 2023; Ravi et al., 2024; Ren et al., 2024a) to segment animal masks in each frame. Next, we apply a state-of-the-art camera estimation algorithm, VGGSfM (Wang et al., 2024a), to optimize the camera’s intrinsic and extrinsic parameters across the video. To ensure a reliable camera estimate, we set thresholds on mean projection errors and mean track lengths, filtering out videos that do not meet these criteria. We then use AnimalAvatar (Sabathier et al., 2024) initialized with Animal3D (Xu et al., 2023) to fit SMAL parameters. Each video is divided into segments of 10 consecutive frames, and AnimalAvatar is applied to each segment independently. This strategy helps mitigate the impact of outliers in camera predictions on the overall optimization quality. To ensure the accuracy of SMAL fitting, we impose thresholds on IoU and PSNR (Hore & Ziou, 2010), filtering out video segments that do not meet our accuracy standards. Once accurate SMAL fittings are obtained, we follow a similar pipeline to extract the desired annotations as used in human cases.
Appendix C Ablation Study
Herein, we conduct additional ablation studies to verify the effectiveness of the proposed designs.
Quantitative Evaluation Regarding Inference Efficiency.
Table 8: Quantitative evaluation regarding inference efficiency. We reduce Stage 1 compute time from 36 seconds to 8 seconds, while maintaining comparable final video quality.
Overall Preference Visual Quality Motion Consistency Amount of Motion
ReVision-SVD (full model) 52.60% 50.27% 51.87% 54.27%
ReVision-SVD (efficient model) 47.40% 49.73% 48.13% 45.73%
Stage 1 can be performed with reduced computational overhead while maintaining similar results. To quantify this, we conducted experiments on $500$ video pairs using the same input images but with varying settings: lower resolution (1:4), fewer frames (8 vs. 16), and fewer denoising steps (32 vs. 50). The efficient model reduced Stage 1 compute time from 36 seconds to 8 seconds. We ran a human preference study on Amazon MTurk, comparing the full and efficient versions for each video, with results provided in Tab. 8. The results demonstrate that the efficient model achieves similar performance to the full model in visual quality and motion consistency, with only a slight drop in the amount of motion, which is likely due to the reduced number of frames. This confirms that Stage 1 can be significantly accelerated with minimal impact on perceived quality.
@@ Appendix C Ablation Study
Parameterized Motion Prior Model. We briefly discuss the quantitative improvements of the proposed PMP in the main paper and provide additional experimental details and results here. To further demonstrate the improvements of PMP, we conducted an additional user study on Amazon MTurk comparing videos generated with and without PMP. Unlike our previous study, which compared our method with SVD, this evaluation focuses on object consistency and motion consistency. We also report the percentage of videos containing incorrect human or animal structures (i.e., the morphological failure rate). We evaluate 500 video pairs, each rated by three different users, resulting in 1,500 total evaluations. The results are presented in Tab. 7 of the main paper. By optimizing with a parametric 3D mesh, our approach significantly reduces incorrect human and animal structures, leading to substantial improvements in object and motion consistency.
Parametric 3D Mesh. Previous human image animation models mainly rely on 2D pose sequences for each frame to provide motion information. However, this approach is not optimal for general video generation. As shown in Fig. 10, we compare the results of using a parametric human mesh model (Loper et al., 2015) versus a human keypoint model (Sun et al., 2019). Our findings indicate that the human mesh model provides a robust object-level prior, which significantly benefits general video generation. Specifically, current video generation models often misinterpret the structure of humans and animals, occasionally producing unrealistic results, such as a man with three arms, an example of morphological failure. This problem becomes more serious in complex motion generation. However, incorporating human and animal priors from 3D mesh models substantially mitigates these structural inaccuracies, enabling more accurate representations of targets.
Figure 10: The parametric 3D mesh serves as an effective object-level prior, ensuring complete human body structures in the coarse video generated during the first stage. In the left two images, the human keypoint model fails to detect the missing right hand, which is accurately “recovered" by the parametric human mesh model. In the right two images, the human mesh model provides a more accurate prior for both blurred hands.
Figure 11: Text prompt and motion strength. We show projected motion sequences generated using different text prompts and motion strength parameters. The text prompt helps define the motion style, while the motion strength controls the speed and complexity of the generated motion. For example, a motion strength of 0.3 results in a simple, direct trajectory, whereas a strength of 0.7 produces a more dynamic and complex motion path. The same target pose is used for the second and third rows to highlight the effect of motion strength.
Text Prompt and Motion Strength. Generating videos from a single image introduces significant ambiguity. To reduce this, we incorporate additional conditioning using a text prompt and a motion strength parameter. Specifically, the text prompt defines the intended motion type, while motion strength controls the speed and complexity of motion within the video. In our experiments, we observe that varying motion strength with the same target pose leads to different motion trajectories. For instance, when moving a hand from point A to B, a video generated with low motion strength results in a direct, simple movement. In contrast, higher motion strength produces a more dynamic and complex trajectory, though it still reaches the same final pose at B. An illustrative example is provided in Fig. 11.
[…]
Model Setting SSIM PSNR LPIPS
VividPose - 0.758 29.83 0.261
Ours (1, 0.5, 0) 0.864 30.08 0.210
Ours (0.8, 0.5, 0.2) 0.851 30.10 0.214
Ours (3, 2, 1) 0.873 30.07 0.217
Ablation on the Confidence Score for Conditioning. We use the confidence map as simple markers to indicate whether a given spatial location contains motion supervision. This is consistent with common practice in prior work (e.g., Emu Video (Girdhar et al., 2023)), where a binary mask is used purely to denote the presence or absence of conditioning signals. In our formulation, assigning different numeric values (e.g., 1 for a full motion sequence, 0.5 for a target pose, and 0 for empty) does not introduce additional learnable meaning. Rather, these values serve as soft indicators that allow us to unify different conditioning types within a single representation, and the model does not depend on their exact magnitudes. We verify this with an ablation in Tab. 9, where we evaluate three parameter sets, i.e., (1, 0.5, 0), (0.8, 0.5, 0.2), and (3, 2, 1), for {full motion sequence, polygon target pose, empty}. The results show that performance is highly robust to the specific choice of confidence values. Given this insensitivity, learning an additional confidence-prediction head is unlikely to provide meaningful gains, while it would introduce extra complexity; therefore, we adopt this simple, standard, and effective marking strategy.
Appendix D Limitations
The proposed method has several remaining limitations. First, it relies on parametric 3D mesh models, requiring multiple off-the-shelf models for different object categories, though it adds only 5 seconds to the total inference time. Recent advances in 3D modeling, such as encoding 3D priors of general objects within a single diffusion model (Liu et al., 2024d), are paving the way for more general, efficient models that can be seamlessly integrated into our pipeline for high-quality video generation. Second, the model still struggles to generate high-quality details such as fingers and hands. Finally, while PMP can generate realistic motion sequences beyond 32 frames, our current implementation, based on vanilla SVD, is limited by memory constraints (80GB RAM). However, recent methods have demonstrated longer video generation ability using pretrained diffusion models (Chen et al., 2023). Exploring long-video generation with 3D knowledge remains future work.
Appendix E Statement of Broader Impact
The proposed ReVision has the potential to facilitate numerous fields through its advanced video generation capabilities. In the realm of creative industries, ReVision can enhance the efficiency and creativity of artists and designers by generating high-fidelity videos. The high-quality generated videos can also contribute to research on synthetic datasets by creating realistic videos, aiding in reducing the annotations required for training vision models. However, with these advancements come ethical considerations, such as the risk of generating deepfakes or other malicious content. It is thus crucial to implement safeguards to minimize potential harms.